A rectangular piece of tin is twice as long as it is wide. A 2 inch square is cut out of each corner, and the sides are turned up to make a box with an open top. The volume of the box is 480 in^3. Find the original dimensions of the piece of tin.
Length:
Width:

Answer by mukhopadhyay(490)

Width = x+4 inch
Length = 2x+8 inch
Height = 2 inch
Volume = (x)(2x+4)(2) = 4x(x+2) = 480
=> x(x+2) = 120
=> x^2+2x-120 = 0
=> (x+12)(x-10) = 0
=> x=10 because it cannot be negative;
Original dimension is (28 inch)X(14 inch)X(2 inch)